\begin{thebibliography}{CDP13}

\bibitem[AD15]{Aumann}
Yonatan Aumann and Yair Dombb.
\newblock The efficiency of fair division with connected pieces.
\newblock {\em ACM Trans. Econ. Comput.}, 3(4):Art. 23, 16, 2015.

\bibitem[Alo87]{Alon}
Noga Alon.
\newblock Splitting necklaces.
\newblock {\em Adv. in Math.}, 63(3):247--253, 1987.

\bibitem[Bar05]{Barbanel}
Julius~B. Barbanel.
\newblock {\em The geometry of efficient fair division}.
\newblock Cambridge University Press, 2005.

\bibitem[BJK13]{BJK}
Steven~J. Brams, Michael~A. Jones, and Christian Klamler.
\newblock \emph{N}-person cake-cutting: There may be no perfect division.
\newblock {\em The American Mathematical Monthly}, 120(1):35--47, 2013.

\bibitem[BT95]{BramsTaylorarticle}
Steven~J. Brams and Alan~D. Taylor.
\newblock An envy-free cake division protocol.
\newblock {\em The American Mathematical Monthly}, 102(1):9--18, 1995.

\bibitem[BT96]{BramsTaylor}
Steven~J. Brams and Alan~D. Taylor.
\newblock {\em Fair division - from cake-cutting to dispute resolution}.
\newblock Cambridge University Press, 1996.

\bibitem[CDP13]{Cech}
Katar{\'{\i}}na Cechl{\'a}rov{\'a}, Jozef Dobo{\v{s}}, and Eva
  Pill{\'a}rov{\'a}.
\newblock On the existence of equitable cake divisions.
\newblock {\em Inform. Sci.}, 228:239--245, 2013.

\bibitem[DS61]{DubinsSpanier}
L.E. Dubins and E.~H. Spanier.
\newblock How to cut a cake fairly.
\newblock {\em The American Mathematical Monthly}, 68(1):1--17, 1961.

\bibitem[Pin76]{Pinkus}
Allan Pinkus.
\newblock A simple proof of the {H}obby-{R}ice theorem.
\newblock {\em Proc. Amer. Math. Soc.}, 60:82--84 (1977), 1976.

\bibitem[Pro16]{Procacciachapter}
Ariel~D. Procaccia.
\newblock Cake cutting algorithms.
\newblock In F.~Brandt, V.~Conitzer, U.~Endriss, J.~Lang, and A.~D. Procaccia,
  editors, {\em Handbook of Computational Social Choice}, chapter~13. Cambridge
  University Press, 2016.

\bibitem[RW97]{RoberstonWebbarticle}
Jack~M. Robertson and William~A. Webb.
\newblock Near exact and envy-free cake division.
\newblock {\em Ars Combinatoria}, 45:97--108, 1997.

\bibitem[RW98]{RobertsonWebb}
Jack~M. Robertson and William~A. Webb.
\newblock {\em Cake-cutting algorithms - be fair if you can}.
\newblock A {K} Peters, 1998.

\bibitem[SS03]{SimmonsSu}
Forest~W. Simmons and Francis~Edward Su.
\newblock Consensus-halving via theorems of {B}orsuk-{U}lam and {T}ucker.
\newblock {\em Math. Social Sci.}, 45(1):15--25, 2003.

\bibitem[Ste48]{Steinhaus}
Hugo Steinhaus.
\newblock {The problem of fair division}.
\newblock {\em Econometrica}, 16(1):101--104, January 1948.

\bibitem[Str80]{Stromquist}
Walter Stromquist.
\newblock How to cut a cake fairly.
\newblock {\em Amer. Math. Monthly}, 87(8):640--644, 1980.

\bibitem[Str81]{Stromquist2}
Walter Stromquist.
\newblock Addendum to: ``{H}ow to cut a cake fairly'' [{A}mer. {M}ath.
  {M}onthly {\bf 87} (1980), no. 8, 640--644;\ {MR} 81m:05016].
\newblock {\em Amer. Math. Monthly}, 88(8):613--614, 1981.

\bibitem[Woo80]{Woodall}
D.R Woodall.
\newblock Dividing a cake fairly.
\newblock {\em Journal of Mathematical Analysis and Applications}, 78(1):233 --
  247, 1980.

\end{thebibliography}

\documentclass{amsart}
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}

\theoremstyle{theorem}
\newtheorem{theorem}{Theorem}

\begin{document}

\title{Existence of a simple and equitable fair division: a short proof
}
\markright{Equitable and simple fair division}
\author[G.~Ch\`eze]{Guillaume Ch\`eze}
\address{Guillaume Ch\`eze: Institut de Math\'ematiques de Toulouse\\
Universit\'e Paul Sabatier Toulouse 3 \\
MIP B\^at. 1R3\\
31 062 TOULOUSE cedex 9, France}
\email{guillaume.cheze@math.univ-toulouse.fr}

\keywords{Fair divisions; Cake cutting; Equitable divisions; Borsuk-Ulam theorem}
\maketitle

\begin{abstract}
In this note we study how to share a good between $n$ players in a simple and equitable way.  We give a short proof for the existence of such fair divisions. \\
\end{abstract}

 In this note we study a fair division problem. Fair divisions problems are sometimes called cake cutting problems. These kinds of problems appear when we study division of land, time or computer memory between different agents with different point of view.  This problem is old: the  ``cut and choose" algorithm already appears in the Bible. In a more scientific way it has been formulated by Steinhaus in 1948, see \cite{Steinhaus}. Nowadays, there exists several articles, see e.g. \cite{DubinsSpanier,BramsTaylorarticle,RoberstonWebbarticle,BJK}, and books, see e.g. \cite{RobertsonWebb,BramsTaylor, Procacciachapter,Barbanel}, about this topic. These results appears in the mathematics, economics, political science, artificial intelligence and computer science literature.\\
 

In this note, our heterogeneous good, e.g. a cake, will be represented by the interval $[0;1]$. 
We consider $n$ players and we associate to each player a non-atomic probability measure $\mu_i$ on the interval $X=[0;1]$ with density $f_i$. These measures represent the utility functions of the player. The set $X$ represents the cake and we want to get a partition of $X=X_1\sqcup \ldots \sqcup X_n$, where the $i$-th player get $X_i$.\\

In this situation several notions of fair division exists:
 \begin{itemize}
 \item Proportional  division: $\forall i,\, \mu_i(X_i) \geq 1/n$.
 \item Exact  division: $\forall i,\, \forall j,\, \mu_i(X_j)=1/n$.
 \item Envy-free division: $\forall i,\, \forall j,\, \mu_i(X_i)\geq \mu_i(X_j)$.
 \item Equitable division: $\forall i,\, \forall j,\, \mu_i(X_i) = \mu_j(X_j)$.\\
 \end{itemize}

  All these fair divisions are possible, see e.g \cite{Steinhaus,DubinsSpanier,BramsTaylor,RobertsonWebb,Cech}. The minimal number of cuts in order to get a fair division has also been studied.\\
  For a proportional fair division $n-1$ cuts are sufficient. This means that for all $i$ there exists an index $j$ such that $X_i= [x_{j},x_{j+1}]$, where $x_0=0$, $x_n=1$. A fair division where each $X_i$ is an interval is called a \emph{simple fair division}. The Banach-Knaster algorithm given in \cite{Steinhaus} shows how to get a simple proportional fair division.\\
   Stromquist and Woodall have shown that we can always get a simple envy-free fair division, see \cite{Stromquist, Stromquist2,Woodall}. These two different proofs use the Brouwer fixed point theorem. \\
  For an exact division $n(n-1)$ cuts are sufficient and this bound is also optimal, see \cite{Alon}. This means that we cannot always get an exact division where each $X_i$ is an interval. 
    The strategy used by Alon to prove his result relies on a general version of the Borsuk-Ulam theorem. \\
    The Borsuk-Ulam theorem is a classical tool in fair division. Indeed, it implies the Ham sandwich theorem see \cite{DubinsSpanier}, and it is also used to get a consensus halving, see e.g. \cite{DubinsSpanier,SimmonsSu}. Consensus halving means that we want to get two sets $A$ and $B$ such that $[0,1]=A \sqcup B$ and for all $i$, $\mu_i(A)=\mu_i(B)$. This result corresponds to the Hobby-Rice theorem and a proof using the Borsuk-Ulam theorem has also been given by Pinkus, see \cite{Pinkus}. As Borsuk-Ulam theorem implies the Brouwer fixed point theorem we can say that the existence of a simple and envy-free fair division relies on the Borsuk-Ulam theorem. This suggests  that the Borsuk-Ulam theorem  is an ubiquitous tool in cake-cutting problems. \\
However  existing proofs about equitable and simple fair division do not use this theorem. Proofs have been given by Cechl\'arov\'a, Dobo\v{s} and Pill\'arov\'a in \cite{Cech} and independently by Aumann and Dombb in \cite{Aumann}.
In \cite{Cech} the authors  have proved that for all permutations $\sigma \in \mathcal{S}_n$ the system of equations:
$$(\star) \quad \int_0^{x_1} f_{\sigma(1)}(x) dx=\int_{x_1}^{x_2} f_{\sigma(2)}(x)dx=\cdots=\int_{x_{n-1}}^1 f_{\sigma(n)}(x)dx,$$
has a solution with $0 \leq x_1 \leq x_2 \leq \cdots \leq x_{n-1} \leq 1$. The permutation $\sigma$ corresponds to the players' order: the $\sigma(i)$-th player get the interval $[x_{i-1},x_i]$.\\
The proof given in \cite{Cech} is based on the notion of the  generalized inverse of a function. 
In \cite{Aumann} the authors give another proof of this result which uses a compactness argument. \\
Thus we have two different proofs but these proofs do not use the classical tool: the Borsuk-Ulam theorem. \\
In the following we show how to get the existence of an equitable and simple fair division  shortly thanks to this classical theorem.

\begin{theorem}
For all densities function $f_i$ and all permutations $\sigma \in \mathcal{S}_n$ the system of equations $(\star)$ has a solution. This means that there exists an equitable and simple fair division. 
\end{theorem}

Our proof relies on the Borsuk-Ulam theorem: If $f:S^k\rightarrow \mathbb{R}^k$ is  continuous and antipodal ($f(-x)=-f(x)$) then there exists $x_0 \in S^k$ such that $f(x_0)=0$.

\begin{proof}
We consider the sphere $S^{n-1}=\{e=(e_1,\dots,e_n) \in \mathbb{R}^{n} | \sum_{i=1}^n e_i^2=1 \}$ and the function
\begin{eqnarray*}
f:S^{n-1} & \rightarrow & \mathbb{R}^{n-1}\\
e&\longmapsto& \big(F_1(e),\ldots,F_{n-1}(e)\big)
\end{eqnarray*}
where 
$$F_i(e)= sgn(e_{i+1}).\int_{e_1^2+\cdots+e_{i}^2}^{e_1^2+\cdots+e_{i}^2+e_{i+1}^2} f_{\sigma(i+1)}(x) dx  - sgn(e_1).\int_{0}^{e_1^2} f_{\sigma(1)}(x) dx.$$
The function $f$ is continuous and antipodal, thus by the Borsuk-Ulam theorem there exists $\tilde{e} \in S$ such that $f(\tilde{e})=0$.\\
We set $x_0=0$ and $x_i=x_{i-1}+\tilde{e}_i^2$ and we get the desired solution for the system $(\star)$.

\end{proof}

\bibliographystyle{alpha} 
 
\bibliography{cakebiblio}

\end{document}